A football is kicked vertically upward from the ground and a student gazing out of a window sees it coming passed him at 16ft/sec. The window is 30ft above the ground.



a. How high does the ball go above the ground?


b. How long does it take to go from a height of 30ft to the highest point?

1
Expert's answer
2021-11-16T10:00:10-0500

"v= 16 \; ft/s \\

h_1 = 30 \;ft \\

g = -32 \;ft/s^2"

a.

"v_f^2 = v_i -2gh \\

0^2 = 16^2 -2 \times 32 \times h_2 \\

h_2 = 4 \;ft"

Height from the ground "= h_1+h_2 = 30 +4 = 34 \;ft"

b.

"v_f =v_i -gt \\

0 = 16 -32t \\

t = \frac{16}{32} = 0.5 \;s"